OpenStudy (anonymous):

What is the value of the expression 7–3? A. –21 B. -1/343 C. 1/ 343 D. 343

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

Are you missing a sign?

OpenStudy (compassionate):

Hi, 7 - 3 = 4. I feel you are missing a sign, as @pabloecortez07 stated.

OpenStudy (anonymous):

7-^3

OpenStudy (compassionate):

Hi, \[7^-3 = -7 * -7 * -7 = -343\] Remember, you cannot have a negative number. Therefore, by putting it over one, we can make it positive. \[\frac{ 1 }{ 343 }\] Thank you for letting me help you! :)

OpenStudy (the_fizicx99):

I guess you can do that but, normally 7^-3 is 1/7^3 So, 7 * 7 * 7 = 343 1/343 :P